OpenStudy (anonymous):

Help please....... A computer can perform c calculations in s seconds. How many minutes will it take the computer to perform k calculations? a. 60ks/c b. ks/c c. ks/60c d. 60c/ks e. k/60cs

OpenStudy (anonymous):

60 sec in 1 min so i think its not b

OpenStudy (anonymous):

i know the answer i want to know how to do it.

OpenStudy (mathstudent55):

The speed of calculations is: \(speed ~of~ calculations = \dfrac{calculations}{time} \) \(speed ~of~calculations = \dfrac{c}{s} \dfrac{calculations}{second} \) \(time = \dfrac{calculations}{speeedd~of~calculations} \) \(\dfrac{1 ~minute}{60 ~seconds} = 1\) \(time = \dfrac{k~calculations}{\frac{c~calculations}{s~seconds}} = \dfrac{ks}{c} seconds \times \dfrac{60~seconds}{1~minute} = \dfrac{60ks}{c} per ~minute\)

OpenStudy (phi):

***A computer can perform c calculations in s seconds. How many minutes will it take the computer to perform k calculations?**** we are given a rate c calcs per s seconds rate*time = amount so \[ \frac{c \ calcs}{s \ secs} \cdot t = k \ calcs\] solve for t \[ t = \frac{k\ s \ secs}{c} \] that is the time in seconds. if we mutliply by 1 minute/60 secs \[ t = \frac{k\ s}{c} \ \cancel{seconds} \cdot \frac{1 \ minute}{60 \ \cancel{seconds}}\\ t= \frac{k s}{60c} \ mins \]

OpenStudy (phi):

mathstudent messed up the conversion from seconds to minutes.

OpenStudy (anonymous):

Is that a formula? The speed of calculations thing he mentioned?

OpenStudy (phi):

rate * time = distance is a formula you learn to figure out how far you go at a given speed. but many problems are similar to this, and can use the same idea when I see amount per time I think "rate" and see if rate * time = amount makes sense. (It usually does)

OpenStudy (anonymous):

ohh okay i see

OpenStudy (anonymous):



OpenStudy (anonymous):

i got that far

OpenStudy (anonymous):

i don't understand from there to be honest

OpenStudy (phi):

what does "x" represent in your ratio?

OpenStudy (anonymous):

# of calculations per minute

OpenStudy (phi):

good. so x is a rate: calculations per minute so use rate * time = amount we know the rate. we are given an amount (k) so we can find time , and it will be in minutes.

OpenStudy (anonymous):

ohhhhhh okay got it

OpenStudy (phi):

it takes practice to think "clearly". Lots of practice.

OpenStudy (anonymous):

yea lots of practice ^_^ thanks phi!!

OpenStudy (mathstudent55):

@phi Thanks for pointing it out. \(time = \dfrac{k~calculations}{\frac{c~calculations}{s~seconds}} = \dfrac{ks}{c} \cancel{seconds} \times \dfrac{1~minute}{60~\cancel{seconds}} = \dfrac{ks}{60c} ~minute\)